OpenStudy (anonymous):

find the slope of the line that passes through each pair of points. (-3,-1) & (-1,5)

hero (hero):

Inserting points (-3,-1) and (-1,5) into y = mx + b results in the following: -1 = -3m + b 5 = -1m + b 3m - 1 = b 1m - 5 = b b = b 3m - 1 = 1m - 5 3m - 1m = -5 + 1 2m = -4 m = -4/2 m = -2 Thus the slope of the line is -2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

False, the slope is 3

OpenStudy (precal):

slope formula